# Convolutional Neural Network NLP Natural Language Processing

### v1

@@ 3.2.1 Datasets
We use three different medical datasets for semantic segmentation. The Japanese Society of Radiological Technology (JSRT) Dataset [54] is publicly available and consists of 247 chest X-rays. Each image comes with landmark annotations [55] for the four anatomical structures of the lungs, heart, and clavicles. For our experiments, we converted the landmarks to dense masks, downsample the dataset to $256\times 256$ pixels, and divide it into a custom split of 160/87 training/test images.
As a second dataset, we include bone segmentation of the paediatric wrist based on the publicly available GRAZPEDWRI-DX (GRAZ) dataset [56]. Here we used human expert segmentation mask of 17 bones provided in [57] of 63 images, which we downsample to $384\times 224$ and split into 43/20 for training/testing. As the last dataset, we include the Tumor InfiltratinG lymphocytes in breast cancER (TIGER) dataset [^3] TIGER training data was accessed on July 25th, 2025 from registry.opendata.aws/tiger., comprising large histology slices. Here, the goal is to segment seven different tissues, including different states/types of tumour and tumour-associated stroma. We selected the subset of slices where the region of interest is provided with dense segmentation mask, obtaining 284 images split into 256/28 training/testing. Since the slices have a very high resolution, we follow a patch-based approach with a patch size of $768\times 768$. During training, we sample one patch per image on-the-fly, with probabilities weighted by the inverse square root of class frequency to oversample minority classes. For inference, we apply a sliding window with 25% overlap and merge predictions using Gaussian weighting, reducing the influence of uncertain patch edges.
@@ 3.2.2 CNN baseline
As a CNN baseline for semantic segmentation, we employ the U-Net [58] with its extension of residual units (MONAI implementation). Each U-Net level employs the same number of channels as the corresponding MetaFormerS12 stage (64, 128, 320, 512). To allow the U-Net to have a sufficiently large receptive field, we deploy an additional fifth stage with 1024 channels. We construct U-Nets with $K\in\{3,5,7\}$ and additionally $K=9$ for the TIGER dataset because of its larger image size. We match the U-Nets’ number of parameters to be as close as possible to the corresponding MetaFormer with convolutional token mixers by varying the number of residual units at each stage. To make sure, each stage is capable of extracting features, we set a minimum of one residual unit.
@@ 3.3 Ranking
To determine the overall best performing token mixer in each setting, we implement a ranking scheme based on the well-established challenges Medical Decathlon [59] and Learn2Reg [60]. Intuitively, every token mixer with a specific kernel size can be considered an individual challenge submission. On each dataset, token mixers are compared against each other and earn a point, for each comparison they significantly won. The results are sorted and token mixers are given scores from the range $[0.1,1]$ per dataset. In cases of ties, they share the average rank score. Finally, an aggregated global rank is computed by the geometric mean across datasets. Since for classification tasks, there is no instance-level performance with the Area under the Receiver Operating Characteristic curve (AUC) metric, we used bootstrapping of the test data with 5k5\,k repeats. With it, we estimate the two-sided 95% confidence interval of the difference in AUC to determine the significance of a token mixers’ comparison. For segmentation, we utilize the Wilcoxon signed-rank test with $p<0.05$ on Dice Similarity Score (DSC).
@@ 4.1.2 Global Ranking of Token Mixers
Grouped convolution with a kernel size of $K=3$ obtains the overall best rank (0.767), followed by pooling with $K=3$ (0.691). To make a more general statement, we averaged the rankings over the different token mixers and kernel sizes (see right column in Tab. 2). We observe that less complex token mixers perform better on average, putting grouped convolution (0.584) first, followed by pooling (0.579), convolution (0.573) and with a large margin local attention (0.268). Moreover, smaller kernel sizes appear to be favourable, with $K=3$ (0.585) and 5 and 7 comparable (0.453/0.466). So in general, less complex token mixers with a smaller kernel sizes perform the best on average across different data sets. See appendix Tab. A.4 for absolute and relative ranking scores for each dataset. The noticeably lower ranks of attention-based token mixers with local on average of 0.268 and global with 0.353 suggest them less suitable when training from scratch on relatively small datasets. However, due to local attention’s training instabilities (see Sec. 4.1.4), we had to reduce the learning rate to $1e^{-4}$. Although, we doubled the number of epochs for a comparable overall optimization budget, the reported AUCs could still be underestimating local attention’s potential. The better average rank of grouped convolution (0.584) than convolution (0.573) suggests that the channel cross-talk lacked by grouped convolution can be effectively compensated by the channel-MLP within the MetaFormer block. Consequently, grouped convolution should be preferred, as it offers lower runtime complexity and reduced parameter count regarding the number of channels (cf. Tab. 1). No token mixing (identity) reaches a competitive place of 4/14 (rank 0.609), which is consistent with the findings of [24]. This raises whether additional token mixing is necessary at all, as the convolution (without any non-linearity) used for patch embedding at the beginning of each stage may already provide sufficient token mixing. Another contributing factor could be the MetaFormer’s extension for global prediction: the final feature map is collapsed via Global Average Pooling (GAP) into a single vector (cf. Fig. 1(c)), which may itself introduce enough token mixing for classification. Furthermore, because each MetaFormer block includes skip connections (cf. Fig. 1(b)), the token mixing by GAP may propagate transitively through all stages of the network.
@@ 4.1.3 Transfer Learning from ImageNet
Another factor that may explain the poor performance of local self-attention, which ranked last in the previous experiment, is its slow convergence and high data requirements as reported in [45]. To mitigate this limitation, we experiment with initializing our models with publicly available ImageNet-pretrained weights. However, since pretrained weights are not available for all our deployed token mixers, we adopt a hybrid MetaFormer variant that uses pooling in the first two stages and global attention in the last two [23]. To avoid a drastic domain shift, we keep pooling as token mixer of the first two stages (introducing no token mixer-parameters) and replace only the global attention in the last two stages with our token mixers. This architecture signature is denoted as $[P,P,T,T]$ (see the corresponding meta row in Tab. 2). Because replacing the token mixer does not modify the channel-MLP, we can leverage pretrained weights for all stages. For self-attention (global and local), we differentiate between two variants: First, transfer the pretrained attention weight matrices $\mathbf{W}_{\{k,v,q,u\}}$ (cf. Sec. 2.2.5) calling it “warm start” and second, initializing their parameters randomly. Since the architecture signatures between the two experiments differs ($[T,T,T,T]$ vs. $[P,P,T,T]$), we also train a $[P,P,T,T]$ variant from scratch (results shown in appendix Tab. A.3), allowing us to measure the impact of pretrained weights.
Fig. 3(a) shows the absolute performance in AUC of the $[P,P,T,T]$ architecture utilizing pretrained weights, and Fig. 3(b) shows the relative improvement when compared to its counterpart trained from scratch. The performance of warm start-local attention on ImageWoof demonstrates that the attention weights of global attention can be effectively reused. As expected, this effect becomes more pronounced as the kernel size of local attention increases, making the two operations more similar. Beyond the benefit of reusing attention weights ($10.54\%,+13.95\%,+15.53\%$), employing pretrained weights for the channel-MLPs alone also yields a substantial improvement ($9.94\%,+11.08\%,+12.13\%$) for local attention trained from scratch. The performance gain is as strong for global attention with ($11.37\%$) and without warm start ($8.29\%$). We also observe performance gains from pre-training across all other token mixers on ImageWoof even if the introduced domain shift diminishes this effect. Notably, the magnitude of improvement correlates with the similarity to the original global self-attention token mixer: for example, parameterized convolutions with the largest kernels possibly benefit more than non-parameterized pooling or no token mixing (identity).
@@
When considering the global ranks, we find 7/8 of the first ranks (18 entries in total, see Tab. 2) holding attention-based token mixers. Local attention with warm start initialization with kernel size $K=7$ and $K=3$ hold the first (0.695) and second (0.613) place, respectively. We attribute the lower ranks of non-attention-based token mixers to the domain gap introduced by replacing the original attention operation, hindering the benefit from pretrained channel-MLPs. However, the relatively low rank (0.526) of global attention with warm-start initialization (8th place) suggests that the superior performance of local attention cannot be explained solely by its similarity to the pretrained global attention, but also by the presence of a local inductive bias. For the non-attention-based token mixers, we observe the same pattern as before, ranking less complex token mixer higher. Grouped convolution and pooling rank similarly, with an average rank of 0.444 and 0.439, followed by convolution (0.405). Interestingly, identity achieves a competitive rank of 0.5426, making it the only non-attention-based token mixer in the top eight. This suggests again that token mixing may be of secondary importance in a global prediction setting, where the final feature map is spatially averaged into a single vector before the classification head. See appendix Tab. A.4 for absolute and relative ranking scores for each dataset.
(a) Area under the Receiver Operating Characteristic curve for the MetaFormer using pretrained ImageNet weights utilizing different token mixers in the last two stages.
(b) Relative improvement from utilizing pretrained weights compared to the variant trained from scratch.
@@ 4.1.4 Training Instabilities of Flex Attention
We employ PyTorch’s flexattention framework to implement the necessary attention pattern for local attention (cf. Eq. (8)). The framework heavily relies on torch.compile to reduce both memory consumption and computational overhead. By monitoring the per-layer gradient $L_{2}$-norm, we observe that compilation can induce gradient explosions within the flexattention module when training on large input resolutions. For classification, this limit is at the spatial resolution of MetaFormer’s first stage (in our setting $56\times 56$). To stabilize training, we experimented with gradient clipping, returning to full 32 bit precision (we usually utilize 16 bit bfloat for efficiency) and also prescaling $\mathbf{q}$ by $\frac{1}{\sqrt{D}}$ before its correlation with $\mathbf{k}$ to reduce the magnitude of values within the matrix multiplication (cf. Eq. 8). In the end, reducing the learning rate from $1e^{-3}$ to $1e^{-4}$ solved the issue for the most situations. To keep the training budget comparable, we double the number of epochs (marked with * in Tab. 2). In cases, where reducing the learning rate was not sufficient, we are forced to disable the compilation of flexattention (marked with $\dagger$ in Tab. 2). This massively increases the resource consumption from one hour training time with around $10\,\text{GB}$ VRAM at a batch size of 128 to 20 hours and $60\,\text{GB}$ at a batch size of 64. To compensate for the reduced batch size, we employ gradient accumulation to simulate an update with the original batch size. Interestingly, we do not encounter any issues when employing the MetaFormer as encoder for segmentation regardless of the larger input sizes up to $768\times 768$ for the TIGER task. We attribute this stability to the SegFormer decoder, which adds shortcuts to each MetaFormer stage (cf. Fig. 1(c)) and thereby shortens the effective gradient path.
@@ 4.2.1 Dataset-specific Performance of Token Mixers
For semantic segmentation as a dense prediction task, we employ the MetaFormer as image encoder with different token mixers at every stage, hence having an architecture signature of $[T,T,T,T]$. We extend its architecture with a lightweight MLP-only decoder of the SegFormer [8] (see Fig. 1(c) and Sec. 3.2 for details). Tab. 3 and Fig. 4 show the average Dice Similarity Score (DSC) across three datasets for different token mixers.
@@ 4.2.2 Global Ranking of Token Mixers
The global rank scores in Tab. 4 (right column) suggest that dense prediction tasks require explicit token mixing at every stage, which was not the case for global prediction tasks (cf. Sec. 4.1.2). We find grouped convolution with a large kernel size $K=7$ yielding the highest global rank (0.832) followed by its $K=5$ variant (0.811). However, we cannot observe the pattern of larger kernels yielding better ranks for the other token mixers. Furthermore, considering the locality of the ranked token mixers (averaged global rank over kernel sizes), from the high locality of grouped/standard convolution (0.746/0.729) to pooling (0.498), local attention (0.294), and finally global attention (0.192), it is evident that a strong local inductive bias is crucial for segmentation in addition to the token mixing itself. This lack of inductive bias hinders the performance of global attention, especially for the small datasets at hand. The identity operation ranks similarly low as global attention does (0.198). As already observed in the classification experiments, using standard convolution (rank 0.729) as the token mixer introduces unnecessary runtime and parameters, since the channel cross-talk that grouped convolution (rank 0.746) would lack is already provided by the MetaFormer block’s channel-MLP. See appendix Tab. A.5 for absolute and relative ranking scores for each dataset.
@@ 5 Conclusion and Future Work
For segmentation tasks, we observe that a strong local inductive bias is crucial, making convolutional token mixers the most effective choice. Here, grouped convolution achieves the highest rank across datasets while offering lower runtime complexity and fewer parameters than standard convolution. We attribute this superior performance to the channel-MLPs, which already provide the channel cross-talk that gets omitted in grouped convolution. Because our segmentation decoder (a lightweight SegFormer design [8]) does not introduce token mixing, this must be provided within the MetaFormer. Consequently, using no token mixing (identity) leads to inferior performance (second to last global rank). We explicitly chose a lightweight decoder to emphasize the impact of different token mixers. Future work might explore alternative decoder designs that incorporate varying degrees of token mixing like MetaSeg [31].
@@
To address the central research question of this work – how to mix tokens (Shaken or Stirred?) within the MetaFormer for medical image processing – our findings suggest that low-complexity token mixers such as grouped convolution, pooling, or even the identity operation are sufficient for classification tasks. We observe no major performance gains from using more complex mixers when training from scratch, confirming that the insights from MetaFormer studies on natural images [24, 23] also generalize to the medical imaging domain. For segmentation tasks, however, we find strong locality crucial and therefore recommend grouped convolution as the preferred token mixer.
@@
Since the medical domain also encompasses volumetric data (e.g., CT or MRI), extending the MetaFormer architecture to 3D inputs represents an important direction for future work. In 3D, the cubically scaling complexity of some token mixers may lead to other recommendations. While our study has focused on classification and segmentation as representative global and dense prediction tasks, other applications such as object detection are also of importance in the medical domain. Future work could explore using the MetaFormer as an image encoder within the Detection Transformer (DETR) framework [9].

### v2

@@ 4.2.1 Datasets
We use three different medical datasets for semantic segmentation. The Japanese Society of Radiological Technology (JSRT) Dataset (Shiraishi et al., 2000) is publicly available and consists of 247 chest X-rays. Each image comes with landmark annotations (Gaggion et al., 2022) for the four anatomical structures of the lungs, heart, and clavicles. For our experiments, we converted the landmarks to dense masks, downsample the dataset to $256\times 256$ pixels, and divide it into a custom split of 160/87 training/test images.
As a second dataset, we include bone segmentation of the paediatric wrist based on the publicly available GRAZPEDWRI-DX (GRAZ) dataset (Nagy et al., 2022). Here we used human expert segmentation mask of 17 bones provided in Keuth et al. (2025) of 63 images, which we downsample to $384\times 224$ and split into 43/20 for training/testing. As the last dataset, we include the Tumor InfiltratinG lymphocytes in breast cancER (TIGER) dataset [^3] TIGER training data was accessed on July 25th, 2025 from registry.opendata.aws/tiger., comprising large histology slices. Here, the goal is to segment seven different tissues, including different states/types of tumour and tumour-associated stroma. We selected the subset of slices where the region of interest is provided with dense segmentation mask, obtaining 284 images split into 256/28 training/testing. Since the slices have a very high resolution, we follow a patch-based approach with a patch size of $768\times 768$. During training, we sample one patch per image on-the-fly, with probabilities weighted by the inverse square root of class frequency to oversample minority classes. For inference, we apply a sliding window with $25\,\%$ overlap and merge predictions using Gaussian weighting, reducing the influence of uncertain patch edges.
@@ 4.2.2 CNN baseline
As a CNN baseline for semantic segmentation, we employ the U-Net (Ronneberger et al., 2015) with its extension of residual units (MONAI implementation). Each U-Net level employs the same number of channels as the corresponding MetaFormerS12 stage (64, 128, 320, 512). To allow the U-Net to have a sufficiently large receptive field, we deploy an additional fifth stage with 1024 channels. We construct U-Nets with $K\in\{3,5,7\}$ and additionally $K=9$ for the TIGER dataset because of its larger image size. We match the U-Nets’ number of parameters to be as close as possible to the corresponding MetaFormer with convolutional token mixers by varying the number of residual units at each stage. To make sure, each stage is capable of extracting features, we set a minimum of one residual unit.
@@ 4.3 Ranking
To determine the overall best performing token mixer in each setting, we implement a ranking scheme based on the well-established challenges Medical Decathlon (Antonelli et al., 2022) and Learn2Reg (Hering et al., 2022). Intuitively, every token mixer with a specific kernel size can be considered an individual challenge submission. On each dataset, token mixers are compared against each other and earn a point, for each comparison they significantly won. The results are sorted and token mixers are given scores from the range $[0.1,1]$ per dataset. In cases of ties, they share the average rank score. Finally, an aggregated global rank is computed by the geometric mean across datasets. Since for classification tasks, there is no instance-level performance with the Area under the Receiver Operating Characteristic curve (AUC) metric, we used bootstrapping of the test data with 5k5\,k repeats. With it, we estimate the two-sided $95\,\%$ confidence interval of the difference in AUC to determine the significance of a token mixers’ comparison. For segmentation, we utilize the Wilcoxon signed-rank test with $p<0.05$ on Dice Similarity Score (DSC).
@@ 5.1.2 Global Ranking of Token Mixers
Grouped convolution with a kernel size of $K=3$ obtains the overall best rank (0.767), followed by pooling with $K=3$ (0.691). To make a more general statement, we averaged the rankings over the different token mixers and kernel sizes (see right column in Tab. 2). We observe that less complex token mixers perform better on average, putting grouped convolution (0.584) first, followed by pooling (0.579), convolution (0.573) and with a large margin local attention (0.268). Moreover, smaller kernel sizes appear to be favourable, with $K=3$ (0.585) and 5 and 7 comparable (0.453/0.466). So in general, less complex token mixers with a smaller kernel sizes perform the best on average across different data sets. See appendix Tab. 7 for absolute and relative ranking scores for each dataset. The noticeably lower ranks of attention-based token mixers with local on average of 0.268 and global with 0.353 suggest them less suitable when training from scratch on relatively small datasets. However, due to local attention’s training instabilities (see Sec. 5.1.4), we had to reduce the learning rate to $1e^{-4}$. Although, we doubled the number of epochs for a comparable overall optimization budget, the reported AUCs could still be underestimating local attention’s potential. The better average rank of grouped convolution (0.584) than convolution (0.573) suggests that the channel cross-talk lacked by grouped convolution can be effectively compensated by the channel-MLP within the MetaFormer block. Consequently, grouped convolution should be preferred, as it offers lower runtime complexity and reduced parameter count regarding the number of channels (cf. Tab. 1). No token mixing (identity) reaches a competitive place of 4/14 (rank 0.609), which is consistent with the findings of Yu et al. (2023). This raises whether additional token mixing is necessary at all, as the convolution (without any non-linearity) used for patch embedding at the beginning of each stage may already provide sufficient token mixing. Another contributing factor could be the MetaFormer’s extension for global prediction: the final feature map is collapsed via Global Average Pooling (GAP) into a single vector (cf. Fig. 2(b)), which may itself introduce enough token mixing for classification. Furthermore, because each MetaFormer block includes skip connections (cf. Fig. 1), the token mixing by GAP may propagate transitively through all stages of the network.
@@ 5.1.3 Transfer Learning from ImageNet
Another factor that may explain the poor performance of local self-attention, which ranked last in the previous experiment, is its slow convergence and high data requirements as reported in Lu et al. (2022). To mitigate this limitation, we experiment with initializing our models with publicly available ImageNet-pretrained weights. However, since pretrained weights are not available for all our deployed token mixers, we adopt a hybrid MetaFormer variant that uses pooling in the first two stages and global attention in the last two (Yu et al., 2022). To avoid a drastic domain shift, we keep pooling as token mixer of the first two stages (introducing no token mixer-parameters) and replace only the global attention in the last two stages with our token mixers. This architecture signature is denoted as $[P,P,T,T]$ (see the corresponding meta row in Tab. 2). Because replacing the token mixer does not modify the channel-MLP, we can leverage pretrained weights for all stages. For self-attention (global and local), we differentiate between two variants: First, transfer the pretrained attention weight matrices $\mathbf{W}_{\{k,v,q,u\}}$ (cf. Sec. 3.2.5) calling it “warm start” and second, initializing their parameters randomly. Since the architecture signatures between the two experiments differs ($[T,T,T,T]$ vs. $[P,P,T,T]$), we also train a $[P,P,T,T]$ variant from scratch (results shown in appendix Tab. 6), allowing us to measure the impact of pretrained weights.
Fig. 4(a) shows the absolute performance in AUC of the $[P,P,T,T]$ architecture utilizing pretrained weights, and Fig. 4(b) shows the relative improvement when compared to its counterpart trained from scratch. The performance of warm start-local attention on ImageWoof demonstrates that the attention weights of global attention can be effectively reused. As expected, this effect becomes more pronounced as the kernel size of local attention increases, making the two operations more similar. Beyond the benefit of reusing attention weights ($10.54\,\%$, $13.95\,\%$, $15.53\,\%$), employing pretrained weights for the channel-MLPs alone also yields a substantial improvement ($9.94\,\%$, $11.08\,\%$, $12.13\,\%$) for local attention trained from scratch. The performance gain is as strong for global attention with ($11.37\,\%$) and without warm start ($8.29\,\%$). We also observe performance gains from pre-training across all other token mixers on ImageWoof even if the introduced domain shift diminishes this effect. Notably, the magnitude of improvement correlates with the similarity to the original global self-attention token mixer: for example, parameterized convolutions with the largest kernels possibly benefit more than non-parameterized pooling or no token mixing (identity).
@@
When considering the global ranks, we find 7/8 of the first ranks (18 entries in total, see Tab. 2) holding attention-based token mixers. Local attention with warm start initialization with kernel size $K=7$ and $K=3$ hold the first (0.695) and second (0.613) place, respectively. We attribute the lower ranks of non-attention-based token mixers to the domain gap introduced by replacing the original attention operation, hindering the benefit from pretrained channel-MLPs. However, the relatively low rank (0.526) of global attention with warm-start initialization (8th place) suggests that the superior performance of local attention cannot be explained solely by its similarity to the pretrained global attention, but also by the presence of a local inductive bias. For the non-attention-based token mixers, we observe the same pattern as before, ranking less complex token mixer higher. Grouped convolution and pooling rank similarly, with an average rank of 0.444 and 0.439, followed by convolution (0.405). Interestingly, identity achieves a competitive rank of 0.5426, making it the only non-attention-based token mixer in the top eight. This suggests again that token mixing may be of secondary importance in a global prediction setting, where the final feature map is spatially averaged into a single vector before the classification head. See appendix Tab. 7 for absolute and relative ranking scores for each dataset.
((a)) Area under the Receiver Operating Characteristic curve for the MetaFormer using pretrained ImageNet weights utilizing different token mixers in the last two stages.
((b)) Relative improvement from utilizing pretrained weights compared to the variant trained from scratch.
@@ 5.1.4 Training Instabilities of Flex Attention
We employ PyTorch’s flexattention framework to implement the necessary attention pattern for local attention (cf. Eq. (8)). The framework heavily relies on torch.compile to reduce both memory consumption and computational overhead. By monitoring the per-layer gradient $L_{2}$-norm, we observe that compilation can induce gradient explosions within the flexattention module when training on large input resolutions. For classification, this limit is at the spatial resolution of MetaFormer’s first stage (in our setting $56\times 56$). To stabilize training, we experimented with gradient clipping, returning to full 32 bit precision (we usually utilize 16 bit bfloat for efficiency) and also prescaling $\mathbf{q}$ by $\frac{1}{\sqrt{D}}$ before its correlation with $\mathbf{k}$ to reduce the magnitude of values within the matrix multiplication (cf. Eq. 8). In the end, reducing the learning rate from $1e^{-3}$ to $1e^{-4}$ solved the issue for the most situations. To keep the training budget comparable, we double the number of epochs (marked with * in Tab. 2). In cases, where reducing the learning rate was not sufficient, we are forced to disable the compilation of flexattention (marked with $\dagger$ in Tab. 2). This massively increases the resource consumption from one hour training time with around $10\,\text{GB}$ VRAM at a batch size of 128 to 20 hours and $60\,\text{GB}$ at a batch size of 64. To compensate for the reduced batch size, we employ gradient accumulation to simulate an update with the original batch size. Interestingly, we do not encounter any issues when employing the MetaFormer as encoder for segmentation regardless of the larger input sizes up to $768\times 768$ for the TIGER task. We attribute this stability to the SegFormer decoder, which adds shortcuts to each MetaFormer stage (cf. Fig. 2(b)) and thereby shortens the effective gradient path.
@@ 5.2.1 Dataset-specific Performance of Token Mixers
For semantic segmentation as a dense prediction task, we employ the MetaFormer as image encoder with different token mixers at every stage, hence having an architecture signature of $[T,T,T,T]$. We extend its architecture with a lightweight MLP-only decoder of the SegFormer (Xie et al., 2021) (see Fig. 2(b) and Sec. 4.2 for details). Tab. 3 and Fig. 5 show the average Dice Similarity Score (DSC) across three datasets for different token mixers.
@@ 5.2.2 Global Ranking of Token Mixers
The global rank scores in Tab. 5 (right column) suggest that dense prediction tasks require explicit token mixing at every stage, which was not the case for global prediction tasks (cf. Sec. 5.1.2). We find grouped convolution with a large kernel size $K=7$ yielding the highest global rank (0.832) followed by its $K=5$ variant (0.811). However, we cannot observe the pattern of larger kernels yielding better ranks for the other token mixers. Furthermore, considering the locality of the ranked token mixers (averaged global rank over kernel sizes), from the high locality of grouped/standard convolution (0.746/0.729) to pooling (0.498), local attention (0.294), and finally global attention (0.192), it is evident that a strong local inductive bias is crucial for segmentation in addition to the token mixing itself. This lack of inductive bias hinders the performance of global attention, especially for the small datasets at hand. The identity operation ranks similarly low as global attention does (0.198). As already observed in the classification experiments, using standard convolution (rank 0.729) as the token mixer introduces unnecessary runtime and parameters, since the channel cross-talk that grouped convolution (rank 0.746) would lack is already provided by the MetaFormer block’s channel-MLP. See appendix Tab. 8 for absolute and relative ranking scores for each dataset.
@@ 6 Conclusion and Future Work
For segmentation tasks, we observe that a strong local inductive bias is crucial, making convolutional token mixers the most effective choice. Here, grouped convolution achieves the highest rank across datasets while offering lower runtime complexity and fewer parameters than standard convolution. We attribute this superior performance to the channel-MLPs, which already provide the channel cross-talk that gets omitted in grouped convolution. Because our segmentation decoder (a lightweight SegFormer design (Xie et al., 2021)) does not introduce token mixing, this must be provided within the MetaFormer. Consequently, using no token mixing (identity) leads to inferior performance (second to last global rank). We explicitly chose a lightweight decoder to emphasize the impact of different token mixers. Future work might explore alternative decoder designs that incorporate varying degrees of token mixing like MetaSeg Kang et al. (2024).
@@
To address the central research question of this work – how to mix tokens (Shaken or Stirred?) within the MetaFormer for medical image processing – our findings suggest that low-complexity token mixers such as grouped convolution, pooling, or even the identity operation are sufficient for classification tasks. We observe no major performance gains from using more complex mixers when training from scratch, confirming that the insights from MetaFormer studies on natural images (Yu et al., 2023, 2022) also generalize to the medical imaging domain. For segmentation tasks, however, we find strong locality crucial and therefore recommend grouped convolution as the preferred token mixer.
@@
Since the medical domain also encompasses volumetric data (e.g., CT or MRI), extending the MetaFormer architecture to 3D inputs represents an important direction for future work. In 3D, the cubically scaling complexity of some token mixers may lead to other recommendations. While our study has focused on classification and segmentation as representative global and dense prediction tasks, other applications such as object detection are also of importance in the medical domain. Future work could explore using the MetaFormer as an image encoder within the Detection Transformer (DETR) framework (Carion et al., 2020).
[…]
This research has been funded by the state of Schleswig-Holstein, Germany, Grant Number 22023005.
[…]
The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals or human subjects.
[…]
We declare we don’t have conflicts of interest.
[…]
All the data used in this work are publicly available and can be accessed through the cited sources. To reproduce the preprocessing steps and our results, please follow the instructions provided in our GitHub repository: https://github.com/multimodallearning/MetaFormerMedImaging/tree/clean_code.
Declaration on Generative AI
During the preparation of this work, the authors used ChatGPT to paraphrase and reword. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the publication’s content.
Appendix A Overview of Appendix
To provide additional insights into our results, we include both accuracy and F1-score as supplementary metrics for the classification tasks. The F1-score might offer more intuition than the AUC, as it represents the harmonic mean of precision and recall, making it a more interpretable indicator of classification performance. Nevertheless, since the AUC evaluates classifier performance across various confidence thresholds, we consider it the primary metric for discussing the classification benchmarks in our work. In addition, we report the absolute and relative ranking scores of all token mixers across datasets for classification and segmentation. The following section provides an overview of the included tables and highlights their aspects that may be of particular interest to the reader.
  • Tab. 4 extends the evaluation of classifiers trained from scratch (MetaFormer with architecture signature $[T$, $T$, $T$, $T]$ and ResNet18) by including accuracy and F1-score. This table provides additional insights into the results discussed in Sec. 5.1.1.
  • Tab. 5 extends the evaluation of classifiers utilizing ImageNet-pretrained weights (MetaFormer with architecture signature $[P,P,T,T]$ and ResNet18) to include accuracy and F1-score. It further complements the discussion presented in Sec. 5.1.3.
  • Tab. 6 reports the AUC and F1-score for the MetaFormer $[P,P,T,T]$ variant trained from scratch. Based on these results and those from Tab. 2 (meta-row $[P,P,T,T]$), the percentage uplift values when compared to pretrained weights are calculated as shown in Fig. 4(b).
  • Tab. 7 presents the absolute and relative ranking scores of different token mixers across classification datasets. This information may be particularly useful for readers seeking the best-performing token mixers for their specific domains.
  • Similar to Tab. 7, Tab. 8 reports the absolute and relative ranking scores of token mixers across segmentation datasets.